Two pipes empty a vat in 1hr and 12 min. When one pipe is used to fill and one to empty the vat, it is filled in 3 hrs. How long would it take each pipe to empty the vat if only one pipe is open?


1
Expert's answer
2022-04-07T12:40:26-0400

x - time of emptying the vat with the first pipe

y - operating time of the second pipe

First pipe alone - (1vat/x minute)(x minute) = 1vat

Second pipe alone - (1vat/y minute)(y minute) = 1vat

Total time 72minute, then 72/x + 72/y = 1 (1)


1/x - amount emptied by the outlet pipe in one hour

1/y - amount filled by the inlot pipe in one hour

1/180 - amount filled by both pipe in one hour (3 hours × 60 = 180 minute)

Then 1/x - 1/y = 1/180 (2)

"\begin{cases}
72/x + 72/y = 1 \\
1/x - 1/y = 1/180
\end{cases}" divide by 72 equation (1)


"\begin{cases}
1/x + 1/y = 1/72 \\
1/x - 1/y = 1/180
\end{cases}"


"2/x = 1/72 + 1/180\\

2/x = 7/360\\

x = 360×2/7\\

x = 103 minute = 1 hour 43 minute"


Substitute x in the first equation

"72/103 + 72/y = 1\\

72/y = 1- 72/103\\

72/y = 31/103\\

y = 72×103/31\\

y = 239 minute = 3 hours 59 minute"


Answer: First pipe - 1 hour 43 minute. Second pipe - 3 hours 59 minute. During this time, each pipe alone will do the work